Glenn Dental Clinic provides general dental care to residents of Philadelphia on a walk-in basis. The clinic has started receiving complaints from patients that the waiting time is too long and has asked you to investigate whether this problem can be solved. Upon arrival, customers first receive a series of paperwork from the receptionist and fill out relevant information such as personal health records and insurance provider. The form is then handed back to a receptionist who enters the information into the computer system for the dentist to see. A dental assistant then takes an X-ray from the patient. A dentist then performs the check-up and discusses any issues with the patient. Based on conversations with staff members at the clinic, you have obtained the following information on the process:

  • It takes about 3 minutes for a customer to fill out the paperwork.  
  • Entry of information on the paperwork into the system and verification with past records takes another 5 minutes for a receptionist. There are two receptionists.
  • It takes 15 minutes on average for the dental assistant to take an X-ray. There are four dental assistants on shift at any moment.       
  • There are ten dentists working at the clinic. Each check-up takes 30 minutes on average.

The following table summarizes the process data collected above.

Resource

Process

Staffing

(Numbers of Workers)

Activity Time

(Minutes per Patient)

Self-Service

Paperwork

-

3

Receptionists

Data Entry

2

5

Dental Assistant

X-ray

4

15

Dentist

Checkup

10

30

Assume that there exists unlimited demand, unless otherwise stated. Assume further that the dentist gets paid $120 per hour while receptionist and dental assistant get paid $30 per hour.

e. Assuming unlimited demand, what would be the utilization at resource “Dentist”? Round to the nearest integer percent. For example, if your answer is 23.456%, input 23; ifyour answer is 23.543%, input 24.

Your answer is .

f. Assume the process started empty, how long (in minutes) would it take 20 patients to go through the system? Round your answer to the nearest integer in minutes. For example, if your answer is 123.45 minutes, input 123.

Your answer is .

g. What is the labor content in minutes? Round your answer to the nearest integer in minutes. For example, if your answer is 123.45 minutes, input 123.

Your answer is .

h. What is the cost of direct labor? Round your answer to the nearest integer in minutes. For example, if your answer is 123.45 minutes, input 123.

Your answer is .
Answer #1

Effective processing time of any activity

= Activity time / Number of staffing

Thus effective processing time at :

Self service = 3/1 = 3

Receptions level = 5/2 = 2.5

Dental assistant X ray = 15/4 = 3.75

Dentist check up = 30/10 = 3

The bottleneck resource will be the one with highest effective time , which in this case is “Dental assistant X ray”

Process cycle time also will be equal to effective processing time of the bottleneck process

Therefore process cycle time = 3.75 minutes/ patient

                 e)Utilization at resource “Dentist”

= Effective processing time at Dentist / Cycle time of the process x 100

= 3/3.75 x 100

= 80%

f)Time it will take for 20 patients to go through the system

= actual processing time for first patient to reach at the doorstep of dentist + Process cycle time per patient x 20 patients

= ( 3 + 5 + 15 ) + 3.75 x 20

= 23 + 75

= 98 minutes

a) Labor content in minutes per patient ( including labor content of the patient for self service at stage I )

= Sum of total time spent at 4 stages

= 3 + 5 + 15 + 30

= 53 minutes

b) Cost per patient

= Sum of ( time taken at each step x wage rate of respective professionals)

Therefore,

Cost of receptionist per patient = 5 minutes x $ 30 per hour/60 minutes = $2.5

Cost of dental assistant per patient = 15 minutes x $30 per hour/60 = $7.5

Cost of dentist per patient = 30 minutes x $120 per hour/60 minutes = $60

Cost of direct labor per patient = Sum of above three = $2.5 +$7.5+$60 = $70